OpenStudy (anonymous):

MEDAL AND FAN! Help me find the zeros of this polynomial function! g(x) = x3 − x2 − 4x + 4

OpenStudy (anonymous):

@earslan100 @enchanted_bubbles @RedNeckOutLaw @TuringTest @IrishBoy123 @pooja195 @sammixboo @dan815 @ganeshie8 @Hero @jagr2713

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

To find the zero of the polynomial function, you need to set it equal to 0 first. g(x) = x^3 − x^2 − 4x + 4 It has a total of three zeros. g(x) = (x-2) (x-1) (x+2) x = 2 x = 1 x = -2

OpenStudy (anonymous):

How did you get to g(x) = (x-2) (x-1) (x+2) please i don't understand how you got there @izuru

OpenStudy (anonymous):

Factor out common terms in the first two terms, then in the last two terms. x^2(x−1)−4(x−1) Factor out the common term x−1 (x−1)(x^2−4) x^2−4 fits the form a^2−b^2 Rewrite it as a = x and b = 2 (x−1)(x^2−22) Use Difference of Squares: a^2−b^2 = (a+b) (a−b) (x−1)(x+2)(x−2)

OpenStudy (anonymous):

OK thank you!